# Compact kinks in a modified Christ-Lee model

### v1

@@ V Kink collisions
Let us consider collisions where a kink or antikink interpolating the vacua $(\phi_{L},\phi_{C})$ moves from the left to the right with velocity $v$, while a kink or antikink interpolating $(\phi_{C},\phi_{R})$ moves from the right to the left with velocity $-v$. We denote this configuration as a $(\phi_{L},\phi_{C})+(\phi_{C},\phi_{R})$ collision, where $\phi_{L}$, $\phi_{C}$, and $\phi_{R}$ are vacua. Before the collision the field can be written exactly as a simple superposition of the moving soliton solutions. When the collision begins, at $t=0$, the field and its time derivative are given by
@@
    Kink-kink collisions $(-1,0)+(0,1)$, which are equivalent to the antikink-antikink collisions $(1,0)+(0,-1)$.
@@
    Kink-antikink collisions $(0,1)+(1,0)$, which are equivalent to the antikink-kink collisions $(0,-1)+(-1,0)$.
@@
    Antikink-kink collisions $(1,0)+(0,1)$, which are equivalent to the kink-antikink collisions $(-1,0)+(0,-1)$.
@@
    Kink-antikink collisions $(-1,1)+(1,-1)$, which are equivalent to the antikink-kink collisions $(1,-1)+(-1,1)$.
@@ V.1 Kink-kink collisions in the $(-1,0)+(0,1)$ sector
We present some simulation results as spacetime color maps for different collision velocities in figures 1416. We plot both the field $\phi(t,x)$ and its energy density $\rho(t,x)$ given by the expression in equation (6). The energy density is plotted in log scale. For the case $n=0.5$ (figure 14) the collision between the kinks is mostly elastic, with no visible emission of radiation. This is expected, since the potential $V_{0.5,0}(\phi)$ has the same shape for $\phi\in[-1,1]$ than the periodic potential studied in ref. [25] for $\phi\in[-2,2]$, apart from a constant scale factor. For that potential, the kink-kink collision was also found to be mostly elastic, with no visible emission of radiation.
Figure 14: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(-1,0)+(0,1)$ collisions with $\epsilon=0$, $n=0.5$, and selected values of $v$.
Figure 15: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(-1,0)+(0,1)$ collisions with $\epsilon=0$, $n=1.5$, and selected values of $v$.
Figure 16: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(-1,0)+(0,1)$ collisions with $\epsilon=0$, $n=3$, and selected values of $v$.
@@ V.2 Kink-antikink collisions in the $(0,1)+(1,0)$ sector
For the case of a kink-antikink collision in the $(0,1)+(1,0)$ sector, the initial conditions are given by
@@
which are even functions of position, i.e. the configuration is symmetric under spatial reflections.
@@
Figure 17: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(0,1)+(1,0)$ collisions with $\epsilon=0$, $n=0.5$, and selected values of $v$.
Figure 18: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(0,1)+(1,0)$ collisions with $\epsilon=0$, $n=1.5$, and selected values of $v$.
Figure 19: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(0,1)+(1,0)$ collisions with $\epsilon=0$, $n=3$, and selected values of $v$.
@@
with $\lambda=1$ for $\phi_{v}=0$, and $\lambda=n$ for $\phi_{v}=\pm 1$. The signum-Gordon equation has both oscillons and shockwaves as known exact solutions [17, 21]. We can introduce new coordinates $x=\lambda^{-1/2}x^{\prime}$, $t=\lambda^{-1/2}t^{\prime}$ to eliminate the $\lambda$ factor from the signum-Gordon equation. Therefore, $\lambda^{-1/2}$ defines a characteristic size for the largest oscillons. In the inner region between the outgoing solitons, the field has values close to the vacuum $\phi_{v}=-1$, therefore the characteristic size for the largest oscillon is $n^{-1/2}$, i.e. it decreases as $n$ increases. This can be observed by visual comparison of figures 1719, specially for the central panel with $v=0.5$. However, since the signum-Gordon equation has scale symmetry, the scale factor will only affect the largest oscillon before higher order effects come into play. This means that smaller oscillons appear in all cases.
@@
Figure 20: Field $\phi(t,0)$ at the middle point $x=0$ between the kink-antikink pair as function of time $t$ and collision velocity $v$ for the $(0,1)+(1,0)$ sector for different values of $n$.
@@
Figure 21: Field $\phi(50,0)$ at the middle point between the kink-antikink pair as function of the collision velocity $v$ and $n$ for the $(0,1)+(1,0)$ sector.
@@ V.3 Antikink-kink collisions in the $(1,0)+(0,1)$ sector
We now study the kink-antikink collisions in the $(1,0)+(0,1)$ sector. The initial conditions are given by
@@
Once again, the configuration is symmetric under spatial reflection. Despite being a qualitatively similar configuration to the $(0,1)+(1,0)$ sector, the $(1,0)+(0,1)$ collision pattern is not equivalent to the former. In particular, the boundary conditions far from the solitons are fundamentally different, approaching the outer minimum of the potential ($\phi=1$) instead of the inner minimum ($\phi=0$). The distinction between the inner and outer minima of the potential is known to produce differences in kink collisions in other models, such as the $\phi^{6}$ [31] and $\phi^{8}$ ones [36].
Figure 22: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(1,0)+(0,1)$ collisions with $\epsilon=0$, $n=0.5$, and some values of $v$.
Figure 23: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(1,0)+(0,1)$ collisions with $\epsilon=0$, $n=1.5$, and some values of $v$.
Figure 24: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for $(1,0)+(0,1)$ collisions with $\epsilon=0$, $n=3$, and some values of $v$.
@@
where $a(t)$ is the $(0,1)$ kink position, we would have $a(0)=x_{0}$ and $a(t)$ would decrease during the initial moments of the collision, only to later increase again. This means that the solitons bounced against each other, instead of passing through as it happened in the $(0,1)+(1,0)$ sector. Another interesting finding is that the formation of central oscillons for high velocities depends on the potential parameter $n$. While a well-defined central oscillon can be seen for $v=0.75$ in the $n=1$ case (figure 22, last column), as $n$ increases the oscillon becomes less clear, while additional radiation starts to dominate (last column of figures 23 and 24).
@@
Figure 25: Field $\phi(t,0)$ at the middle point between the antikink-kink pair as function of time $t$ and collision velocity $v$ for the $(1,0)+(0,1)$ sector with different value of $n$.
@@
Figure 26: Field $\phi(50,0)$ at the middle point between the antikink-kink pair as function of the collision velocity $v$ and $n$ for the $(1,0)+(0,1)$ sector.
@@ V.4 Kink-antikink collisions in the $(-1,1)+(1,-1)$ sector
For the last case, concerning the $(-1,1)+(1,-1)$ kink-antikink collisions, the initial conditions are
@@
We first look at the case $\epsilon=0$, for which the potential has a triple well shape. For this case, the kinks $(-1,1)$ are simply a superposition of a $(-1,0)$ and $(0,1)$ kinks with zero separation and no overlap. Since each subkink is compact, they do not interact and can therefore be separated simply by acquiring different velocities. Hence, a $(-1,1)+(1,-1)$ collision for $\epsilon=0$ is actually a four soliton collision $(-1,0)+(0,1)+(1,0)+(0,-1)$.
Figure 27: Field $\phi(t,x)$ for $(-1,1)+(1,-1)$ collisions with $\epsilon=0$ and selected values of $n$ and $v$.
We present some examples of the simulation results for $\epsilon=0$ in figure 27. The presence of four solitons generates a highly non-trivial dynamics, with several possible outcomes. For small velocities, we observe cases with a complete annihilation of the kink-antikink pair, such as the cases $n=0.5$ with $v=0.15$, $n=1$ with $v=0.3$, and $n=3.0$ with $v=0.1$, displayed in the first column of the plots. There are also cases in which only the inner subkinks $(0,1)$ and $(1,0)$ annihilate, but the outer subkinks $(-1,0)$ and $(0,-1)$ are able to escape. These include cases with a formation of a large central oscillon ($n=0.5$ with $v=0.75$, $n=1$ with $v=0.62$, and $n=3$ with $v=0.2$, for example), as well as cases without it ($n=1$ with $v=0.5$ and $n=3$ with $v=0.7$, for example). The case $n=3$ with $v=0.7$ is specially interesting, since it features two large oscillons that accompany the outgoing solitons. We also observe cases in which all four solitons escape, such as $n=0.5$ with $v=0.6$, $n=1$ with $v=0.8$, and $n=3$ with $v=0.9$, for example. In the first two of these examples, the $(-1,1)$ kink and the $(1,-1)$ antikink are broken into their composing subkinks, i.e. each one of the subkinks and sub-antikinks escapes with a different velocity, forming two growing regions where $\phi=0$ (dark regions in the plots). In the last example, $n=3$ with $v=0.9$, the $(-1,1)$ kink and $(1,-1)$ antikink remain intact.
Figure 28: Field $\phi(t,0)$ for $(-1,1)+(1,-1)$ collisions as a function of the velocity $v$ for selected values of $n$ in the case $\epsilon=0$, for which the potential has a triple well shape.
@@
Figure 29: Field $\phi(50,0)$ for $(-1,1)+(1,-1)$ collisions as function of velocity $v$ and $n$ in the case $\epsilon=0$, for which the potential has a triple well shape.
The dependence on $n$ can be more systematically explored by looking at the field at the origin on later times. Choosing $t=50$, in figure 29, we present the values of $\phi(50,0)$ as function of $n$ and $v$. Looking at the lower region of the plot, we see that the formation of a second capture window is a feature mostly of values of $n$ only slightly larger than $0.5$. In general, we see that the capture window grows with $n$, specially for partial captures.
Figure 30: Field $\phi(t,x)$ for $(-1,1)+(1,-1)$ collisions with $\epsilon=0.1$ and selected values of $n$ and $v$.
Proceeding further, let us look at the collisions in the case $\epsilon=0.1$, for which the potential features a false vacuum. Some examples of simulations for this case are presented in figure 30. As discussed previously, for the false vacuum cases, the $(-1,1)$ kink can still be interpreted as being mostly composed of two subkinks. The main new feature present in these simulations is the formation of false vacuum bubbles after the annihilation of the inner subkink-antisubkink pair. These bubbles grow in size as the outer subkinks escape. However, since the $\phi=0$ vacuum is false, the bubble requires energy to grow. At a certain point, the false vacuum bubble stops growing and shrinks back, with outer subkinks moving towards one another and eventually colliding. After the subkink collisions, there are several possible outcomes. The entire structure can be annihilated and radiate away ($n=3$ with $v=0.5$, for example), a new bubble can form ($n=0.5$ with $v=0.4$, for example), or the original $(-1,1)+(1,-1)$ kink-antikink pair can reemerge and escape ($n=0.5$ with $v=0.6$, for example).
Figure 31: Field $\phi(t,0)$ for $(-1,1)+(1,-1)$ collisions as function of velocity $v$ for selected values of $n$ in the case $\epsilon=0.1$, for which the potential has a false vacuum.
By looking at the field at the origin as a function of $v$ and $t$ in figure 31, we see that the reappearance of the $(-1,1)+(1,-1)$ is present only in the $n=0.5$ case for the time interval simulated ($0\leq t\leq 50$). The bubbles in $n=0.5$ are also special by mostly appearing together with a large central oscillon. For other values of $n$, we also observe large central oscillons, but bubbles without them are also common (black regions in figure 31).
Figure 32: Field $\phi(t,x)$ for $(-1,1)+(1,-1)$ collisions with $\epsilon=\infty$ and selected values of $n$ and $v$.
@@
Figure 33: Field $\phi(t,0)$ for $(-1,1)+(1,-1)$ collisions as function of velocity $v$ for selected values of $n$ in the case $\epsilon=\infty$, for which the potential has a double well shape.
Figure 34: Field $\phi(50,0)$ for $(-1,1)+(1,-1)$ collisions as function of velocity $v$ and $n$ in the case $\epsilon=\infty$, for which the potential has a double well shape.
We may also analyse the dependence on $n$ by considering the late time field $\phi(50,0)$ in figure 34. Looking at the lower region of the plot, we see that only for $n$ slightly larger than $0.5$ we have capture for all velocities. For larger values of $n$, there is a critical velocity separating the escape cases, as well as one or two windows of bounces followed by escape.
Figure 35: Field $\phi(50,0)$ for $(-1,1)+(1,-1)$ collisions as function of velocity $v$ and $\epsilon$ for selected values of $n$.

### v2

@@ V Kink collisions
[…]
@@
    Kink-kink collisions (1,0)+(0,1)(-1,0)+(0,1)( - 1 , 0 ) + ( 0 , 1 ), which are equivalent to the antikink-antikink collisions (1,0)+(0,1)(1,0)+(0,-1)( 1 , 0 ) + ( 0 , - 1 ).
@@
    Kink-antikink collisions (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ), which are equivalent to the antikink-kink collisions (0,1)+(1,0)(0,-1)+(-1,0)( 0 , - 1 ) + ( - 1 , 0 ).
@@
    Antikink-kink collisions (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ), which are equivalent to the kink-antikink collisions (1,0)+(0,1)(-1,0)+(0,-1)( - 1 , 0 ) + ( 0 , - 1 ).
@@
    Kink-antikink collisions $+(1,-1)$ ( 1 , - 1 ), which are equivalent to the antikink-kink collisions $+(-1,1)$ ( - 1 , 1 ).
@@ V.1 Kink-kink collisions in the $(-1,0)+(0,1)$ sector
We present some simulation results as spacetime color maps for different collision velocities in figures 1416. We plot both the field $\phi(t,x)$ and its energy density $\rho(t,x)$ given by the expression in equation (6). The energy density is plotted in log scale. For the case $n=0.5$ (figure 14) the collision between the kinks is mostly elastic, with no visible emission of radiation. This is expected, since the potential $V_{0.5,0}(\phi)$ has the same shape for $\phi\in[-1,1]$ than the periodic potential studied in ref. [25] for $\phi\in[-2,2]$, apart from a constant scale factor. For that potential, the kink-kink collision is also mostly elastic, with no visible emission of radiation.
Figure 14: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (1,0)+(0,1)(-1,0)+(0,1)( - 1 , 0 ) + ( 0 , 1 ) collisions with $\epsilon=0$, $n=0.5$, and selected values of $v$.
Figure 15: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (1,0)+(0,1)(-1,0)+(0,1)( - 1 , 0 ) + ( 0 , 1 ) collisions with $\epsilon=0$, $n=1.5$, and selected values of $v$.
Figure 16: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (1,0)+(0,1)(-1,0)+(0,1)( - 1 , 0 ) + ( 0 , 1 ) collisions with $\epsilon=0$, $n=3$, and selected values of $v$.
@@ V.2 Kink-antikink collisions in the $(0,1)+(1,0)$ sector
For the case of a kink-antikink collision in the (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) sector, the initial conditions are given by
@@
which are even functions of position, i.e., the configuration is symmetric under spatial reflection.
@@
Figure 17: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) collisions with $\epsilon=0$, $n=0.5$, and selected values of $v$.
Figure 18: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) collisions with $\epsilon=0$, $n=1.5$, and selected values of $v$.
Figure 19: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) collisions with $\epsilon=0$, $n=3$, and selected values of $v$.
@@
with $\lambda=1$ for $\phi_{v}=0$, and $\lambda=n$ for $\phi_{v}=\pm 1$. The signum-Gordon equation has both oscillons and shockwaves as known exact solutions [17, 21]. We can introduce new coordinates $x=\lambda^{-1/2}x^{\prime}$, $t=\lambda^{-1/2}t^{\prime}$ to eliminate the $\lambda$ factor from the signum-Gordon equation. Therefore, $\lambda^{-1/2}$ defines a characteristic size for the largest oscillons. In the inner region between the outgoing solitons, the field has values close to the vacuum $\phi_{v}=-1$, therefore the characteristic size for the largest oscillon is $n^{-1/2}$, i.e., it decreases as $n$ increases. This can be observed by visual comparison of figures 1719, specially for the central panel with $v=0.5$. However, since the signum-Gordon equation has scale symmetry, the scale factor will only affect the largest oscillon before higher order effects come into play. This means that smaller oscillons appear in all cases.
@@
Figure 20: Field $\phi(t,0)$ at the middle point $x=0$ between the kink-antikink pair as function of time $t$ and collision velocity $v$ for the (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) sector for different values of $n$.
@@
Figure 21: Field $\phi(50,0)$ at the middle point between the kink-antikink pair as function of the collision velocity $v$ and $n$ for the (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) sector.
@@ V.3 Antikink-kink collisions in the $(1,0)+(0,1)$ sector
We now study the kink-antikink collisions in the (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) sector. The initial conditions are given by
@@
Once again, the configuration is symmetric under spatial reflection. Despite being a qualitatively similar configuration to the (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) sector, the (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) collision pattern is not equivalent to the former. In particular, the boundary conditions far from the solitons are fundamentally different, approaching the outer minimum of the potential ($\phi=1$) instead of the inner minimum ($\phi=0$). The distinction between the inner and outer minima of the potential is known to produce differences in kink collisions in other models, such as the $\phi^{6}$ [27] and $\phi^{8}$ ones [30].
Figure 22: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) collisions with $\epsilon=0$, $n=0.5$, and some values of $v$.
Figure 23: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) collisions with $\epsilon=0$, $n=1.5$, and some values of $v$.
Figure 24: Field $\phi(t,x)$ and energy density $\rho(t,x)$ for (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) collisions with $\epsilon=0$, $n=3$, and some values of $v$.
@@
where $a(t)$ is the (0,1)(0,1)( 0 , 1 ) kink position, we would have $a(0)=x_{0}$ and $a(t)$ would decrease during the initial moments of the collision, only to later increase again. This means that the solitons bounced against each other, instead of passing through as it happened in the (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) sector. Another interesting finding is that the formation of central oscillons for high velocities depends on the potential parameter $n$. While a well-defined central oscillon can be seen for $v=0.75$ in the $n=1$ case (figure 22, last column), as $n$ increases the oscillon becomes less clear, while additional radiation starts to dominate (last column of figures 23 and 24).
@@
Figure 25: Field $\phi(t,0)$ at the middle point between the antikink-kink pair as function of time $t$ and collision velocity $v$ for the (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) sector with different value of $n$.
@@
Figure 26: Field $\phi(50,0)$ at the middle point between the antikink-kink pair as function of the collision velocity $v$ and $n$ for the (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) sector.
@@ V.4 Kink-antikink collisions in the $(-1,1)+(1,-1)$ sector
For the last case, concerning the $+(1,-1)$ ( 1 , - 1 ) kink-antikink collisions, the initial conditions are
@@
We first look at the case $\epsilon=0$, for which the potential has a triple well shape. For this case, the kinks (1,1)(-1,1)( - 1 , 1 ) are simply a superposition of a (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) kinks with zero separation and no overlap. Since each subkink is compact, they do not interact and can therefore be separated simply by acquiring different velocities. Hence, a $+(1,-1)$ ( 1 , - 1 ) collision for $\epsilon=0$ is actually a four soliton collision $+(1,0)+(0,-1)$ ( 1 , 0 ) + ( 0 , - 1 ).
Figure 27: Field $\phi(t,x)$ for $+(1,-1)$ ( 1 , - 1 ) collisions with $\epsilon=0$ and selected values of $n$ and $v$.
We present some examples of the simulation results for $\epsilon=0$ in figure 27. The presence of four solitons generates a highly non-trivial dynamics, with several possible outcomes. For small velocities, we observe cases with a complete annihilation of the kink-antikink pair, such as the cases $n=0.5$ with $v=0.15$, $n=1$ with $v=0.3$, and $n=3.0$ with $v=0.1$, displayed in the first column of the plots. There are also cases in which only the inner subkinks (0,1)(0,1)( 0 , 1 ) and (1,0)(1,0)( 1 , 0 ) annihilate, but the outer subkinks (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,-1)( 0 , - 1 ) are able to escape. These include cases with a formation of a large central oscillon ($n=0.5$ with $v=0.75$, $n=1$ with $v=0.62$, and $n=3$ with $v=0.2$, for example), as well as cases without it ($n=1$ with $v=0.5$ and $n=3$ with $v=0.7$, for example). The case $n=3$ with $v=0.7$ is specially interesting, since it features two large oscillons that accompany the outgoing solitons. We also observe cases in which all four solitons escape, such as $n=0.5$ with $v=0.6$, $n=1$ with $v=0.8$, and $n=3$ with $v=0.9$, for example. In the first two of these examples, the (1,1)(-1,1)( - 1 , 1 ) kink and the (1,1)(1,-1)( 1 , - 1 ) antikink are broken into their composing subkinks, i.e., each one of the subkinks and sub-antikinks escapes with a different velocity, forming two growing regions where $\phi=0$ (dark regions in the plots). In the last example, $n=3$ with $v=0.9$, the (1,1)(-1,1)( - 1 , 1 ) kink and (1,1)(1,-1)( 1 , - 1 ) antikink remain intact.
Figure 28: Field $\phi(t,0)$ for $+(1,-1)$ ( 1 , - 1 ) collisions as a function of the velocity $v$ for selected values of $n$ in the case $\epsilon=0$, for which the potential has a triple well shape.
@@
Figure 29: Field $\phi(50,0)$ for $+(1,-1)$ ( 1 , - 1 ) collisions as function of velocity $v$ and $n$ in the case $\epsilon=0$, for which the potential has a triple well shape.
The dependence on $n$ can be more systematically explored by looking at the field at the origin on later times. Choosing $t=50$, in figure 29, we present the values of $\phi(50,0)$ as function of $n$ and $v$. Looking at the lower region of the plot, we see that the formation of a second capture window is a feature mostly of values of $n$ only slightly larger than 0.50.50.5. In general, we see that the capture window grows with $n$, specially for partial captures.
Figure 30: Field $\phi(t,x)$ for $+(1,-1)$ ( 1 , - 1 ) collisions with $\epsilon=0.1$ and selected values of $n$ and $v$.
Proceeding further, let us look at the collisions in the case $\epsilon=0.1$, for which the potential features a false vacuum. Some examples of simulations for this case are presented in figure 30. As discussed previously, for the false vacuum cases, the (1,1)(-1,1)( - 1 , 1 ) kink can still be interpreted as being mostly composed of two subkinks. The main new feature present in these simulations is the formation of false vacuum bubbles after the annihilation of the inner subkink-antisubkink pair. These bubbles grow in size as the outer subkinks escape. However, since the $\phi=0$ vacuum is false, the bubble requires energy to grow. At a certain point, the false vacuum bubble stops growing and shrinks back, with outer subkinks moving towards one another and eventually colliding. After the subkink collisions, there are several possible outcomes. The entire structure can be annihilated and radiate away ($n=3$ with $v=0.5$, for example), a new bubble can form ($n=0.5$ with $v=0.4$, for example), or the original $+(1,-1)$ ( 1 , - 1 ) kink-antikink pair can reemerge and escape ($n=0.5$ with $v=0.6$, for example).
Figure 31: Field $\phi(t,0)$ for $+(1,-1)$ ( 1 , - 1 ) collisions as function of velocity $v$ for selected values of $n$ in the case $\epsilon=0.1$, for which the potential has a false vacuum.
By looking at the field at the origin as a function of $v$ and $t$ in figure 31, we see that the reappearance of the $+(1,-1)$ ( 1 , - 1 ) is present only in the $n=0.5$ case for the time interval simulated ($0\leq t\leq 50$). The bubbles in $n=0.5$ are also special by mostly appearing together with a large central oscillon. For other values of $n$, we also observe large central oscillons, but bubbles without them are also common (black regions in figure 31).
Figure 32: Field $\phi(t,x)$ for $+(1,-1)$ ( 1 , - 1 ) collisions with $\epsilon=\infty$ and selected values of $n$ and $v$.
@@
Figure 33: Field $\phi(t,0)$ for $+(1,-1)$ ( 1 , - 1 ) collisions as function of velocity $v$ for selected values of $n$ in the case $\epsilon=\infty$, for which the potential has a double well shape.
Figure 34: Field $\phi(50,0)$ for $+(1,-1)$ ( 1 , - 1 ) collisions as function of velocity $v$ and $n$ in the case $\epsilon=\infty$, for which the potential has a double well shape.
We may also analyse the dependence on $n$ by considering the late time field $\phi(50,0)$ in figure 34. Looking at the lower region of the plot, we see that only for $n$ slightly larger than 0.50.50.5 we have capture for all velocities. For larger values of $n$, there is a critical velocity separating the escape cases, as well as one or two windows of bounces followed by escape.
Figure 35: Field $\phi(50,0)$ for $+(1,-1)$ ( 1 , - 1 ) collisions as function of velocity $v$ and $\epsilon$ for selected values of $n$.